The cockroach Periplaneta americana can detect a static electric field of magnitude 6.00 kN/C using their long antennae. If the excess static charge on a cockroach is modeled as point charges located at the end of each antenna, what magnitude of charge q would each antenna possess in order for each antennae to experience a force of magnitude 4.50 ?N from the external electric field? Calculate q in units of nanocoulombs (nC) .
Answer #1

QE = 4.5×10^-6

Q = (4.5×10^-6) / (6×10^3) = 0.75nC

Q: charge on the antennae of cockroach

E: static electric field

formula used : Force = QE